OpenStudy (anonymous):

question about break even point. I am not to sure how to tackle this, so any help would be greatly appreciated. "find the break even point of each pair of equation. round to nearest decimal answers to the nearest tenth." E=26 +3.2x I=4.5x im not sure what E and i stand for.

OpenStudy (unklerhaukus):

is there something in the question you haven't included?

OpenStudy (anonymous):

ohhhhh im sorry it's E=26 + 3.2x sorry, sort of new here.

OpenStudy (unklerhaukus):

i think the break even point is where the lines cross

OpenStudy (unklerhaukus):

ie, the point that solves both the equations

OpenStudy (anonymous):

so how would i go about this? should i just graph it in my calculator or is there some formula for this?

OpenStudy (unklerhaukus):

well i wold simply solve the second equation for x (by dividing by 4.5) and then substituting this into the first equation

OpenStudy (anonymous):

so im looking for i?

OpenStudy (unklerhaukus):

your looking for the \(x\) that makes I=E

OpenStudy (anonymous):

i got x= i/4.5 then plugged it into 26+ 3.2 (i/4.5) am i on the right path?

OpenStudy (unklerhaukus):

yeah, E=26+3.2 (i/4.5) simplify this a bit and you will have E in terms of I

OpenStudy (anonymous):

okay, so im stuck at E= 26 +3.2i/4,5 i dont think i can those two values.

OpenStudy (anonymous):

add those*

OpenStudy (unklerhaukus):

actually i think there is an better way to do this problem The break even point is when I=E so solve 26 +3.2x=4.5x

OpenStudy (anonymous):

ohhh okay.

OpenStudy (unklerhaukus):

can you find \(x\) now?

OpenStudy (anonymous):

Yes thank you, you are too awesome. i got x=20

OpenStudy (unklerhaukus):

good stuff, can you round that to the nearest tenth?

OpenStudy (anonymous):

do i need to? it broke evenly.

OpenStudy (unklerhaukus):

im not sure,

OpenStudy (anonymous):

i think the instructions might be in case it didnt break down evenly.

OpenStudy (unklerhaukus):

To be a point the point must have an \(x\) and a \(y\) value you've found \(x=20\) , to find \(y\) substitute \(x\) into either equation to get \(y\), then write the point like this \((x,y)\)

OpenStudy (anonymous):

Y=i =4.5 * 20= 90 (20,90)?

OpenStudy (unklerhaukus):

Thats it \[(20,90)\checkmark\]

OpenStudy (anonymous):

awesome, possum :) thanks for the help mate.